Required information [The following information applies to the questions displayed below.] Preble Company manufactures one product. Its variable manufacturing overhead is applied to production based on direct labor-hours and its standard cost card per unit is as follows: Direct material: 5 pounds at $8.00 per pound $ 40.00 Direct labor: 2 hours at $14 per hour 28.00 Variable overhead: 2 hours at $5 per hour 10.00 Total standard variable cost per unit $ 78.00 The company also established the following cost formulas for its selling expenses: Fixed Cost per Month Variable Cost per Unit Sold Advertising $ 200,000 Sales salaries and commissions $ 100,000 $ 12.00 Shipping expenses $ 3.00 The planning budget for March was based on producing and selling 25,000 units. However, during March the company actually produced and sold 30,000 units and incurred the following costs: Purchased 160,000 pounds of raw materials at a cost of $7.50 per pound. All of this material was used in production. Direct-laborers worked 55,000 hours at a rate of $15.00 per hour. Total variable manufacturing overhead for the month was $280,500. Total advertising, sales salaries and commissions, and shipping expenses were $210,000, $455,000, and $115,000, respectively.

Required: 1. What raw materials cost would be included in the company’s flexible budget for March?

2. What is the materials quantity variance for March? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.). Input the amount as a positive value.)

3. What is the materials price variance for March? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.). Input the amount as a positive value.)
Answer #1

1) Raw material in Flexible budget = 30000*5*8 = $1200000

2) Material quantity variance = (30000*5-160000)*8= 80000 U

3) Material price variance = (8-7.50)*160000 = 80000 F